OpenStudy (anonymous):

Write the quadratic equation in vertex form... y=(2/3)x^2+4x-1?

OpenStudy (anonymous):

multiply it in 3/2 ------ > (3/2)y=x^2+6x-(3/2) ---> (3/2)y+ (3/2)+9 =x^2+6x+9

OpenStudy (anonymous):

(3/2)y+(21/2)=(x+3)^2

OpenStudy (anonymous):

(3/2)(y+7)=(x+3)^2

OpenStudy (anonymous):

vertex = (-3,-7)

OpenStudy (anonymous):

that's it.